OpenStudy (anonymous):

the new problem is write the percent as decimal. round to the nearest thousandth if the divistion does not temiate 4/32%

OpenStudy (anonymous):

i posted my next problem 23 minutes ago on the left side of the screen but no one has helped me yet.

OpenStudy (anonymous):

32% = 32/100 = 0.32 However, since this does not include a thousandth value I figure the question must have been different, are you sure you have posted it right? Perhaps it should be, (4/32)%=4/32=0.125 %

OpenStudy (anonymous):

if this is as you have written it i.e. 4/32% then divide 4/(32/100) to give answer, or as I got above, 4/0.32 = 12.5